The base of an isosceles triangle is 11 in. less than the length of the equal sides. If the perimeter of the triangle is 37 in., find the length of each of the sides.
Answer by rajagopalan(174)

Let the length of each the equal sides = x
The base of an isosceles triangle is 11 in. less than that = x-11
Side 1 =x
Side2=x
Side3=x-11
Add all=perimeter=3x-11
Perimeter of the triangle is 37 in
So the length of each of the sides is calculated like this:
3x-11=37
3x=37+11
3x=48
x=48/3
x=16
okay
good day